# Depth Matters: Exploring Deep Interactions of RGB-D for Semantic Segmentation in Traffic Scenes

### v1

@@ I INTRODUCTION
With the rapid development of autonomous driving, the ability to accurately understand complex traffic scenes has become increasingly crucial. Traditional single-sensor data often fail short under varying lighting conditions and with dynamic objects. To overcome these challenges, combining texture information with spatial geometry, RGB-D based apporaches [1, 2, 3] have energed as the mainstream methods.
However, existing methods often treat depth information as auxiliary and overlook the intrinsic properties of depth maps. As a result, conflicts between RGB and depth map in sensitive regions lead to misalignment in the learned attention maps, significantly affecting perception in complex real-world scenes. For instance, the attention shift shown in Fig. 1(a, b) leads to substantial prediction errors.
@@ III-C Depth Linear Cross Attention
The attenntion effectively combines the global information from the RGB features with the local information from the depth features. In this formula, the dot product $Q\cdot K^{T}$ between the query matrix $Q$ and the key matrix $K$ is computed for explore the similarity of RGB-D feature spaces. The scale factor $scale$ is employed to balance the influence of features from different modalities. The similarity attention is then normalized using the $softmax$ to generate the attention weight matrix. The weights are subsequently applied to the value matrix $V$ to achieve the sufficient fusion of RGB-D feature spaces. It is denoted as:
@@ III-E Complexity
In Depth LCA, adaptive average pooling reduces the feature dimension from $H\times W\times C$ to $P\times P\times C$. Consequently, the computational complexity of Depth LCA decreases from $O(HWC)$ to $O(P^{2}C)$, where $P$ is much smaller than $H$ and $W$. Moreover, the computational complexity of MLP Decoder is $O(HWC_{in}M)$, which is lower compared to the traditional comvolutional decoder’s complexity of $O(HWC^{2})$, where $C_{in}$ and $M$ are much smaller than $C$. The reduced computational complexity ensures that our design is lightweight and efficient, effectively meeting the real-time demands of traffic scene understanding.
@@ IV EXPERIMENTS
IV-A Datasets and Metric
IV-B Implementation Details
IV-C Comprehensive Evaluation
@@ V CONCLUSIONS
Limitations. One notable limitation of our work is still relies on the quality of the depth iamges to some extent. Despite our design mitigates error accumulation, the improvement contributed by sparse depth is not significant, as shown in Fig. 1(f: Cityscapes). Therefore, we will explore an unified framework that integrates segmentation with depth estimation to reduce reliance on depth images and strengthen the relationship between pixels and depth.
@@ ACKNOWLEDGMENT
The authors would like to thank the editor and the anonymous reviewers for their critical and constructive comments and suggestions.

### v2

@@ I INTRODUCTION
With the rapid development of autonomous driving, the ability to accurately understand complex traffic scenes has become increasingly crucial. Traditional single-sensor data often fail short under varying lighting conditions and with dynamic objects. To overcome these challenges, combining texture information with spatial geometry, RGB-D based approaches [1, 2, 3] have emerged as the mainstream methods.
However, existing methods often treat depth information as auxiliary and overlook the intrinsic properties of depth maps[4]. As a result, conflicts between RGB and depth map in sensitive regions lead to misalignment in the learned attention maps, significantly affecting perception in complex real-world scenes. For instance, the attention shift shown in Fig. 1(a, b) leads to substantial prediction errors.
@@ III-C Depth Linear Cross Attention
The attenntion effectively combines the global information from the RGB features with the local information from the depth features. In this formula, the dot product $Q\cdot K^{T}$ between the query matrix $Q$ and the key matrix $K$ is computed to explore the similarity of RGB-D feature spaces. The scale factor $scale$ is employed to balance the influence of features from different modalities. The similarity attention is then normalized using the $softmax$ to generate the attention weight matrix. The weights are subsequently applied to the value matrix $V$ to achieve the sufficient fusion of RGB-D feature spaces. It is denoted as:
@@ III-E Complexity
In Depth LCA, adaptive average pooling reduces the feature dimension from $H\times W\times C$ to $P\times P\times C$. Consequently, the computational complexity of Depth LCA decreases from $O(HWC)$ to $O(P^{2}C)$, where $P$ is much smaller than $H$ and $W$. Moreover, the computational complexity of MLP Decoder is $O(HWC_{in}M)$, which is lower compared to the traditional convolutional decoder’s complexity of $O(HWC^{2})$, where $C_{in}$ and $M$ are much smaller than $C$. The reduced computational complexity ensures that our design is lightweight and efficient, effectively meeting the real-time demands of traffic scene understanding.
@@ IV EXPERIMENTS
IV-A Datasets and Metric
IV-B Implementation Details
IV-C Comprehensive Evaluation
@@ V CONCLUSIONS
Limitations. One notable limitation of our work is still relies on the quality of the depth images to some extent. Despite our design mitigates error accumulation, the improvement contributed by sparse depth is not significant, as shown in Fig. 1(f: Cityscapes). Therefore, we will explore an unified framework that integrates segmentation with depth estimation to reduce reliance on depth images and strengthen the relationship between pixels and depth.
@@ ACKNOWLEDGMENT
This work was supported in part by the Natural Science Foundation of Xiamen, China, under Grant 3502Z202373036; in part by the National Natural Science Foundation of China under Grant 42371457, Grant 42301468, and Grant 42371343; in part by the Key Project of Natural Science Foundation of Fujian Province, China, under Grant 2022J02045; in part by the Natural Science Foundation of Fujian Province, China, under Grant 2022J01337, Grant 2022J01819, Grant 2023J01801, Grant 2023J01799, Grant 2022J05157, and Grant 2022J011394.